I tried using the sto key for this on my calculator but I don't think I put it in properly. I tried different ways to substitute the answers with the questions but I couldn't find the solution. The question is:
'Maria's 7-hour drive to college was 355 miles long. She averaged 45 miles per hour for the first 2 hours. Which of the following is closest to her average speed in miles per hour, for the remainder of her drive?
A.) 38
B.) 51
C.) 53
D.) 55
E.) 56

Found 2 solutions by rothauserc, ikleyn:
Answer by rothauserc(4718)

time = 7 - 2 = 5 hours
:
distance = 355 - (2 * 45) = 265
:
*******************************************************
average speed for remainder of drive = 265 / 5 = 53mph
:
answer is C
*******************************************************
:

Answer by ikleyn(52906)

.
The trick is to apply your brain/mind first.